Katy invests in a total of $24,000 in two accounts. The first account earned a rate of return of 12% ( after a year). However, in the second account suffered a 8% loss in the same period. At the end of one year , the total amount of money gained was $-120.00. How much was invested into each account?
Answer by VFBundy(438)

12% return:
Principal = p
Rate = 0.12
Interest = 0.12p

8% loss:
Principal = 24000 - p
Rate = -0.08
Interest = -0.08(24000 - p) = -1920 + 0.08p

0.12p + (-1920 + 0.08p) = -120

0.12p - 1920 + 0.08p = -120

0.20p = 1800

p = 9000

12% return:
Principal = p = $9000

8% loss:
Principal = 24000 - p = $15000